OpenStudy (anonymous):

y = 3x + 4 2x + 3y = 10 Which of the following steps could be used to solve by substitution? 3x + 4 = 2x + 3y 2x + 3(3x + 4) = 10 2(3x + 4) + 3y = 10 3(y = 3x + 4)

jimthompson5910 (jim_thompson5910):

Hint: \[ \Large {\color{red}{y}} = 3x+4\\ \Large 2x + 3{\color{red}{y}} = 10 \]

OpenStudy (anonymous):

@jim_thompson5910 i dont get the hint..

jimthompson5910 (jim_thompson5910):

the two 'y's are the same, so you can replace the 3y in the second equation with 3(...) the dot dot dot is whatever y is equivalent to

jimthompson5910 (jim_thompson5910):

example \[ \Large \color{red}{y} = x+2\\ \Large 5x + 10\color{red}{y} = 17 \] After substitution, we get \[ \Large 5x + 10\color{red}{y} = 17\\ \Large 5x + 10(\color{red}{y}) = 17\\ \Large 5x + 10(\color{red}{x+2}) = 17 \]

OpenStudy (anonymous):

@jim_thompson5910 could you possibly tell me what you got so i can see how you got it

jimthompson5910 (jim_thompson5910):

do you see in my example how I replaced y with x+2 ?

OpenStudy (anonymous):

@jim_thompson5910 i suck @ math and this my last answer on my homework

jimthompson5910 (jim_thompson5910):

all you're doing is replacing y with what it is equivalent to. Look back at my example again

jimthompson5910 (jim_thompson5910):

In my example, y is equivalent to x+2 so that's how I replaced the y with x+2 in the second equation. Don't forget the parenthesis

OpenStudy (study_together):

It's B because \[2x+3(3x+4)=10\]can be substituted to \[5x+7=10\]and this is the substitution property.

OpenStudy (marihelenh):

The first equation shows what y is equal to. \[y=3x+4\] So, in the second equation, you can put what the y was equal to in the place of where the y was. \[2x+3y=10\] \[2x+3(3x+4)=10\]